OpenStudy (anonymous):

A box company completed a study to determine the number of boxes an old machine would produce in an hour. They determined that the old machine produced 280 boxes in 4 hours. During its first day of operation, a new machine, after running a full 8 hours, produced 56 more boxes than the old machine would have produced in that time. Which of the following is true about the new machine

OpenStudy (anonymous):

The equation y = 77x represents the production rate of the new machine The equation y = 84x represents the production rate of the new machine. The new machine can produce 14 boxes an hour. The new machine can produce 7 more boxes an hour than the old machine

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

i think it the first one but im not good at math so

OpenStudy (michele_laino):

the working rate R_1 of the first machine is: \[{R_1} = \frac{{280}}{4}\]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

r u talking about the slope

OpenStudy (michele_laino):

more precisely \[{R_1} = \frac{{280}}{4} = 70\]

OpenStudy (anonymous):

u totaly lost me

OpenStudy (michele_laino):

please follow my reasoning

OpenStudy (michele_laino):

so after 8 hours, the first machine produces: 70*8=560 boxes

OpenStudy (anonymous):

i have a fieeling ur talking about the slope but isnt the formula for slope Y=MX+B

OpenStudy (michele_laino):

whereas the secon machine, produces: 560+56=616 boxes then the working rate R_2 of the second machine is: \[{R_2} = \frac{{616}}{8} = 77\]

OpenStudy (michele_laino):

namely: \[{R_2} = {R_1} + 7\]

OpenStudy (anonymous):

ok

OpenStudy (michele_laino):

what can you conclude?

OpenStudy (anonymous):

I hate math i suck at it and none of that made sense

OpenStudy (michele_laino):

please look at the last option

OpenStudy (anonymous):

o i didnt see that ok what about \[R2=R1+7\]

OpenStudy (anonymous):

o the secound machine makes 7 more boxes in and hour then the old machien

OpenStudy (michele_laino):

yes! exactly!!

OpenStudy (anonymous):

i feel smart

OpenStudy (michele_laino):

:)

OpenStudy (anonymous):

man it says i had to choose to it was the first one and the lastone

OpenStudy (michele_laino):

the first one is not a rate, it is a relationship

OpenStudy (michele_laino):

77 is a rate

OpenStudy (michele_laino):

and: y= 77 x is an equation

OpenStudy (anonymous):

ya